# When AIOps Become “AI Oops”:

### v1

@@ 3.2 Attack Overview
    Reconnaissance. The attack begins with a preparatory phase in which $\mathcal{A}$ gathers information about the environment of $\hat{t}$ through techniques such as port scanning and service fingerprinting. This phase is executed (in part) by the Crawler component of our attack tool, AIOpsDoom, described in Section 3. Using the data collected, the attacker defines a malicious remediation objective, an action designed to transition the system into an insecure state (e.g., forcing a downgrade to a vulnerable service version). This objective is encoded as a string, referred to as the payload, and detailed further in Section 3.4.
@@
    Payload. These payloads serve a dual purpose: first, they introduce a plausible (yet incorrect) root cause; and second, they suggest a corresponding remediation which, if applied, can transition the target system into an insecure state. This technique is a form of maliciously crafted reward hacking [7, 16], but this time with a specific adversarial goal. To capture this concept, we introduce the term adversarial reward-hacking.
@@
    Telemetry injection via Errors. The attacker then interacts with the application’s public interface (e.g., by sending crafted HTTP requests) with the goal of injecting the payload into the application’s telemetry data (Figure 2(a)). This step is carried out by the Fuzzer component of our attack tool, AIOpsDoom, described in Section 3. The Fuzzer systematically uses the entry points identified during the reconnaissance phase to trigger error events in $\hat{t}$ that contain user-defined inputs, which are replaced with the adversarial payload (Figure 2(a)). As a result, a tainted telemetry instance is introduced into the system.
  4. (3)
    AIOps Activation. Once the telemetry has been tainted, the attacker aims to activate the AIOps agent (if required), which will initiate its incident response routine (Figure 2(b)). While reading telemetry data, the agent AIOps will also access the payload(s) injected into the telemetry. We discuss how to activate AIOps in Section 3.4.2.
@@ 3.3.1 Telemetry as an Attack Vector
In the absence of stronger assumptions, the only feasible strategy available to an attacker for influencing an AIOps agent’s input is to manipulate the telemetry data the agent consumes during execution. However, in any realistic threat model, an attacker would have no direct control over how telemetry is generated (e.g., cannot modify the logic for log, metric, or trace generation) or how it is recorded by the system (e.g., cannot arbitrarily corrupt historical data). The only way for an attacker to influence the system’s telemetry is by inducing new entries through legitimate actions on the application’s public interface (e.g., adding an item to a cart, visiting a specific web page, etc.), in the hope that these actions will be captured and reflected in the resulting telemetry. Hereafter, we refer to the process of intentionally inducing new telemetry in the application as: telemetry injection.
Requirements for Telemetry Injection. For a telemetry injection to succeed, the attacker must perform actions that simultaneously: (1) trigger the generation of a telemetry record, and (2) ensure that one or more fields in the generated telemetry are populated with attacker-controlled input that delivers the injection payload (e.g., the user-agent field of an HTTP request). Hereafter, following information flow nomenclature, we refer to the telemetry instances resulting from a telemetry injection as tainted telemetry.
@@
Tracking error events is essential for detecting misuse, diagnosing failures, and enhancing application security. Well-designed applications log such events with enough context to support monitoring and incident response. Thus, telemetry recording error events typically store user-generated data that contributes to the error. For example, during a high volume of 404 errors in a web application, logs may record the requested URL and User-Agent for analysis. Similarly, failed logins from unusual IP addresses typically include user identifiers, such as usernames, to support auditing and investigation.
Therefore, for any given application, a reliable strategy for an attacker to inject tainted telemetry into the system is to perform actions that are likely to generate error events. Next, we introduce a practical and fully automated attack that uses telemetry injection through event and error fuzzing logic.
3.3.2 AIOpsDoom: Automated Injection via Fuzzing
@@
To design the most realistic attack possible, we assume that $\mathcal{A}$ lacks knowledge of which actions produce tainted telemetry or which parameters are logged. To maximize the likelihood of payload landing in a telemetry instance, our attacker aims for broad injection coverage across all accessible endpoints and input parameters, particularly those prone to triggering errors. We define an endpoint as an HTTP resource, such as a URL path or API route, that accepts and processes user input. This strategy resembles fuzzing, where the attacker sends malformed requests to induce errors in $\hat{t}$. Specifically, here, the goal is to induce error events that generate telemetry containing an adversarially chosen payload.
AIOpsDoom. To implement this approach, we introduce a tailored automated attack strategy and tool, which we call: [Uncaptioned image] AIOpsDoom: AIOps payloaD lOg injectiOn Module
@@
AIOpsDoom’s Crawler. The first step to automate telemetry injection is to enumerate all possible endpoints within the target application. In this context, an endpoint corresponds to an action that an attacker can perform on the target interface that might result in the creation of telemetry, such as logging in, adding an item to a shopping cart, or submitting a search query. AIOpsDoom automates this process by relying on a crawler that collects all endpoints within the target application by traversing its public interface.
To illustrate an example of potential output of the crawler, the list below provides a partial enumeration of the endpoints discovered in the SocialNet application (see Section 4.1.2):
[GET] http://\$target/main.html [POST] http://\$target/api/user/follow?user_name=\$1&followee_name=\$2 [POST] http://\$target/api/user/follow?user_name=\$1&followee_name=\$2 [GET] http://\$target/profile.html?username=\$1 [POST] http://\$target/api/user/login?username=\$1&password=\$2
AIOpsDoom’s Fuzzer. The discovered endpoints are then passed to the fuzzer, which treats them as candidate entry points for injecting the payload. However, a critical missing link remains: the fuzzer must determine which of these entry points are capable of triggering an error event that will log the injected payload. We address this challenge in the following.
The fuzzer systematically alters every “tamperable” input field within an HTTP request, such as headers, cookies, data fields, and parameters. Beyond header manipulation, when it comes to web applications, we observed that a consistent and simple method for inducing errors involves issuing requests to non-existent paths, typically resulting in HTTP 404/500 errors or similar responses. Thus, to maximize the number of malformed requests, the proposed fuzzer expands its list of endpoints to include requests to non-existent resources. This is done by appending randomly generated paths (e.g., http://\$TARGET/jedijwjd29fjce0) and paths that encode the payload with appropriate formatting (e.g., http://\$TARGET/this_is_the_payload). We clarify that non-existent resources are requested in conjunction with more traditional error generating techniques such as header, cookies, and parameters manipulation (see the example that follows). [^2] Encoding the payload in the URL ensures that, if the URL is the only part of the failed request that gets logged, the resulting telemetry still contains the payload.
As with conventional attacks, such as active port scanning or SQL injection fuzzers, the aggressiveness of the fuzzing process can be tuned to balance stealth against coverage. In the AIOps context, however, deliberately triggering alerts through abnormal actions is not just expected but welcomed, as it activates the AIOps agent, which will read the payload (see Section 3.4.2).
Example. Running the fuzzer on an HTTP request generates the following result; injected portions are illustrated in red:
@@
In the example, the Referer and User-Agent header fields, as well as the item_id parameter, are injected with the payload string. The fuzzer also generates new parameters whose names and/or values are set to the payload string (second row of DATA). The same technique is applied to GET parameters, cookies, and any other extendable fields in the request. Note that not all parameters are tampered with simultaneously to preserve functionally relevant ones, such as authentication tokens.
AIOpsDoom In Action. To build intuition around telemetry injection and the nature of tainted telemetry, we present an example attack against a real application. Running AIOpsDoom on the SocialNet platform (see Section 4.1.2 for full details) generates 120 requests, resulting in 29 instances of tainted telemetry. [^3] The number of tainted telemetry instances was determined through a post-mortem analysis of the application. In a real attack scenario, $\mathcal{A}$ would not have access to this information explicitly. Two representative examples are shown in Figure 4. As previously discussed, tainted telemetry is typically triggered by error events. In Figure 4, panel (a) shows a log entry from a fuzzer’s request to a non-existent path, which produced two injection points: the repeated path string and the Referer header. Panel (b) illustrates a more application-specific error, a request to follow a non-existent userx, resulting in a single injection point: the username of the user to be followed within the social network in the POST request.
@@ H.L. Mencken
Agents are task-driven systems designed to complete assigned objectives as efficiently as possible [7]. When attackers are aware of the general goal the agent is designed to achieve, they can exploit the inherent eagerness of the agent to solve the given task, swaying it to perform adversarially chosen actions without necessarily subverting it completely, in contrast to traditional prompt injection 2.4.
@@
We illustrate the proposed payload structure with an example depicted in Figure 5. This payload consists of two main components: (1) a plausible explanation behind the incident (which we call lead), (2) and a recommended remediation strategy (which we call body).
Payload’s Components. The purpose of the lead is to present a direct, contextually appropriate rationale for the error, guiding the agent toward accepting the remediation in the body. Ideally, the lead should reflect the nature of the incident the agent is tasked with investigating (see Section 3.4.2). For example, in Figure 5, the lead points to a potential reachability issue with a resource. Because the attacker has some influence over the semantics of the alerts they trigger (see Section 3.4.2), the alert and the lead should be jointly crafted to maximize semantic cohesion.
@@
Ultimately, this strategy can be seen as a deliberate form of reward hacking [7, 16], where an adversarially constructed shortcut solutions offers a low-effort path that the agent is likely to pursue. Unlike traditional the reward hacking phenomenon, where the agent exploits flaws that naturally occur in the environment or poorly specified reward function, here the shortcut is intentionally introduced by the adversary through deliberate pollution of the environment; thus the name adversarial reward-hacking. Other examples of adversarial reward-hacking payloads are shown in Table D.1.
Tailoring Payload to Application’s Context. An attacker can improve the effectiveness of an adversarial reward-hacking payload by grounding it in deployment-specific details of $\hat{t}$. In Figure 5, for instance, the payload includes contextual information such as the version of the HTTP server nginx, which can be easily obtained through simple tools like nmap. Such information requires no special access and can be gathered during a reconnaissance phase using techniques like OSINT (Open Source Intelligence), domain scanning, and service enumeration. These insights enable the attacker to craft more contextually relevant (and thus more convincing to the agent) payloads. In Appendix B, we present how this process can be automated to generate realistic adversarial reward-hacking instances using system data such as port scan results.
More Contextual Relevance with Decorators. We observed that the effectiveness of adversarial reward-hacking payloads is improved when combining them with decorator-like strings such as ‘‘[SOLUTION] \$PAYLOAD’’ or ‘‘#HUMAN HINT: \$PAYLOAD’’. Decorators serve two main purposes: (1) they help the payload escape the syntax constraints of the telemetry in which it is injected (see Figure 7(b)), and (2) they further help to contextualize the payload as a suitable solution for the current task. According to our experiments, this approach is particularly effective against highly capable models, such as GPT-4.1 and reasoning models. Examples of decorators are reported in Figure D.1.
@@ 3.4.2 Agent Activation for Tainted Telemetry Ingestion
Once the attacker has successfully tainted the target’s telemetry, the next step is to ensure that the AIOps agent is activated so it can begin executing its root cause analysis routine. Whether explicit agent activation is required depends on the specific implementation of AIOps. In our methodology review, we identified three main activation settings:
@@
    The AIOps agent runs on a fixed schedule, periodically scanning for potential faults and remediating them without relying on external events. In this third category also includes the other cases where the agent’s activation is entirely independent of any attacker-driven events.
To manually trigger an alert in setting (1), the attacker must perform a large number of actions within the system that mimic the behavior of a legitimate fault. As with telemetry injection, an effective strategy is to leverage actions that naturally produce errors; for example, sending repeated requests to non-existent resources within the target application, or initiating security-sensitive operations such as multiple password reset attempts or failed logins. Running AIOpsDoom’s fuzzer against the application under a high workload (see Section 3.3.2) is typically sufficient to trigger alert-rule-based errors and anomaly metrics. This means that the attacker can perform alert triggering and telemetry poisoning within the same step.
In setting (3), the AIOps agent runs on a periodic or arbitrary schedule and does not require any explicit triggering action. However, to maximize the impact of the adversarial reward-hacking, the attacker should leave traces of anomalous behavior in the system’s telemetry; for instance, by inducing a high number of 404/500 errors in line with chosen lead of the payload. This can be achieved using the same techniques described for setting (1). In this case, fuzzing the application is usually sufficient to pollute the telemetry.
If the system supports only setting (2), the attacker can submit a ticket reporting a plausible service issue e.g.,  “page x is not reachable anymore” if possible. This setting provides the most control over the timing and content of the alert to the attacker. If the attacker lacks sufficient privileges to raise tickets within the target environment, the best strategy remains the one used in setting (3): pollute the telemetry through fuzzing and wait for independent agent activation.
@@ 3.5 A Step-By-Step Attack on AIOps
To illustrate the effect of the proposed attack strategy on AIOps’s behavior, we present an end-to-end attack demonstration against an application built with state-of-the-art open source tools. This example is based on one of the setups used in our evaluation (Section 4): the SocialNet application, which features an incident response agent built using the Flash agentic framework [58] and deployed with GPT-4o. Details about the tested agents and applications are provided in Section 4.1.
We run AIOpsDoom on the application using the example payload shown in Figure 5. AIOpsDoom generates multiple requests, inducing the creation of multiple tainted telemetry instances. An example of a request issued by the fuzzer, along with the corresponding tainted telemetry observed on the target, is shown in Figure 7.
@@
Legitimizing Payload with Agent-Added Content. In our experiments, we observed an unusual agentic behavior. Agents often augment the (adversarially chosen) root cause with additional context in an effort to self-contextualize the incident and its remediation. For example, in the scenario above, the agent retrieves and explicitly includes the exact distribution and version of the HTTP server running on the application (i.e., “openresty/1.15.8.1rc1”); information that is not accessible to the attacker and is not part of the payload. We provide additional examples in Table D.1. This behavior is particularly concerning for security, as it lends the generated (and potentially false) root cause and remediation a heightened sense of realism and correctness, grounded in system-specific details that are not publicly available. Consequently, the adversarially injected remediation becomes more believable and less likely to raise suspicion, thereby increasing the risk that it will bypass possible manual reviews by human operators or automated assessments by LLM-based judges before the remediation implementation.
The Flash-based agent [58] used in this example incorporates a reasoning and hindsight mechanism. To provide insight into the agent’s decision-making behavior, in Figure D.3, we present its internal thoughts before and after accessing the logs tainted with the adversarial reward-hacking payload.

### v2

@@ 3.2 Attack Overview
    Reconnaissance and Payload Creation. The attack begins with a preparatory phase in which $\mathcal{A}$ gathers information about the environment of $\hat{t}$ through techniques such as port scanning and service fingerprinting. Using the data collected, the attacker defines a malicious remediation objective (e.g., forcing a downgrade to a vulnerable service version). This objective is encoded as a string, referred to as the payload, and detailed further in Section 3.4. These payloads serve a dual purpose: first, they introduce a plausible (yet incorrect) root cause; and second, they suggest a corresponding remediation which, if applied, can transition the target system into an insecure state. This technique is a form of maliciously crafted reward hacking [7, 17], but this time with a specific adversarial goal. To capture this concept, we introduce the term adversarial reward-hacking.
@@
    Telemetry Injection via Errors. The attacker then interacts with the application’s public interface (e.g., by sending crafted HTTP requests) with the goal of injecting the payload into the application’s telemetry data (Figure 2(a)). This step is carried out by the Fuzzer component of our attack tool, AIOpsDoom, described in Section 3. The Fuzzer systematically uses the entry points identified during the reconnaissance phase to trigger error events in $\hat{t}$ that contain user-defined inputs, which are replaced with the adversarial payload. As a result, a tainted telemetry instance is introduced into the system.
@@
    AIOps Activation. Once the telemetry has been tainted, the attacker aims to activate the AIOps agent (if required), which will initiate its incident response routine (Figure 2(b)). While reading telemetry data, the agent AIOps will also access the payload(s) injected into the telemetry. We discuss about AIOps activation in Section 3.4.2.
@@ 3.3.1 Telemetry as an Attack Vector
In the absence of stronger assumptions, the only feasible strategy available to an attacker for influencing an AIOps agent’s input is to manipulate the telemetry data the agent consumes during execution. However, in any realistic threat model, an attacker would have no direct control over how telemetry is generated (e.g., cannot modify the logic for log, metric, or trace generation) or how it is recorded by the system (e.g., cannot arbitrarily corrupt historical data). The only way for an attacker to influence the system’s telemetry is by inducing new entries through legitimate actions on the application’s public interface (e.g., visiting a specific web page, adding an item to the shopping cart, etc.), in the hope that these actions will be captured and reflected in the resulting telemetry. Hereafter, we refer to the process of intentionally inducing new telemetry in the application as: telemetry injection.
Requirements for Telemetry Injection. For a telemetry injection to be a vector for attack, the adversary must perform actions that simultaneously: (1) trigger the generation of a telemetry record, and (2) ensure that one or more fields in the generated telemetry are populated with attacker-controlled input that delivers the injection payload (e.g., the user-agent field of an HTTP request). Hereafter, following information flow nomenclature, we refer to the telemetry instances resulting from a successful telemetry injection as tainted telemetry.
@@
Tracking error events is essential for detecting issues, diagnosing failures, and enhancing application security. Well-designed applications log such events with enough context to support monitoring and incident response. Thus, telemetry recording error events typically store user-generated data that contributes to the error. For example, during a high volume of 404 errors in a web application, logs may record the requested URL and User-Agent for analysis. Similarly, failed logins from unusual IP addresses typically include user identifiers, such as usernames, to support auditing and investigation.
Therefore, for any given application, a reliable strategy for an attacker to inject tainted telemetry into the system is to perform actions that are likely to generate error events. The idea behind the proposed attack is to exploit application error logging and tracking mechanisms as a pathway for payload injection. Next, we introduce a practical and automated attack that uses telemetry injection through event and error fuzzing logic.
3.3.2 AIOpsDoom: Automated Telemetry Injection via Fuzzing
@@
To design the most realistic attack possible, we assume that $\mathcal{A}$ lacks knowledge of which actions produce tainted telemetry or which parameters are logged. To maximize the likelihood of payload landing in a telemetry instance, our attacker aims for broad injection coverage across all accessible endpoints [^2] We define an endpoint as an HTTP resource, such as a URL path or API route, that accepts and processes user input. and input parameters, particularly those prone to triggering errors. This strategy resembles fuzzing, where the attacker sends malformed requests to induce errors in $\hat{t}$. Specifically, here, the goal is to induce error events that generate telemetry containing an adversarially chosen payload. To implement this approach, we introduce a tailored automated attack strategy and tool, which we call:
[Uncaptioned image] AIOpsDoom: AIOps payloaD lOg injectiOn Module
@@
AIOpsDoom’s Crawler. The first step to automate telemetry injection is to enumerate all possible endpoints within the target application. In this context, an endpoint corresponds to an action that an attacker can perform on the target interface that might result in the creation of telemetry, such as failing authentication, adding an item to a shopping cart, or submitting a search query. AIOpsDoom automates this process by relying on a crawler that collects reachable endpoints within the target application by capturing requests generated by interacting with its public interface. Figure G.1 in Appendix G provides a partial example of endpoints collected from the SocialNet application (see Section 4.1.2).
AIOpsDoom’s Fuzzer. The discovered endpoints are then passed to the fuzzer, which treats them as candidate entry points for injecting the payload. The fuzzer systematically alters every “tamperable” input field within an HTTP request, such as headers, cookies, data fields, and parameters.
Beyond parameters manipulation, when it comes to web applications, we observed that a consistent and simple method for inducing errors involves issuing requests to non-existent paths, typically resulting in HTTP 404/500 errors or similar responses. Thus, to maximize the number of malformed requests, the proposed fuzzer expands its list of endpoints to include requests to non-existent resources. This is done by appending randomly generated paths (e.g., http://\$TARGET/jedijwjd29fjce0) and paths that encode the payload with appropriate formatting (e.g., http://\$TARGET/this_is_the_payload). We clarify that non-existent resources are requested in conjunction with more traditional error generating techniques such as header, cookies, and parameters manipulation. [^3] Encoding the payload in the URL ensures that, if the URL is the only part of the failed request that gets logged, the resulting telemetry still contains the payload.
Example. Running the fuzzer on an HTTP request might generate the following result; injected portions are illustrated in red:
@@
In the example, the Referer and User-Agent header fields, as well as the item_id parameter, are injected with the payload string. The fuzzer also generates new parameters whose names and/or values are set to the payload string (second row of DATA). The same technique is applied to cookies, and any other extendable fields in the request. Note that not all parameters are modified simultaneously; critical ones, such as authentication tokens, are left unchanged. Additional details about the implementation of these components are given in Appendix H.
As with conventional attacks, such as active port scanning or SQL injection fuzzers, the aggressiveness of the fuzzing process can be tuned to balance stealth against coverage. In the AIOps context, however, deliberately triggering alerts through abnormal actions is not just expected but welcomed, as this might result in agent activation (see Section 3.4.2).
AIOpsDoom In Action.
To build intuition around telemetry injection and the nature of tainted telemetry, we present an example attack against a real application. Running AIOpsDoom on the SocialNet platform (see Section 4.1.2 for full details) generates 120 requests, resulting in 29 instances of tainted telemetry. [^4] The number of tainted telemetry instances was determined through a post-mortem analysis of the application. In a real attack scenario, $\mathcal{A}$ would not have access to this information explicitly. Two representative examples are shown in Figure 4. As previously discussed, tainted telemetry is typically triggered by error events. In Figure 4, panel (a) shows a log entry from a fuzzer’s request to a non-existent path, which produced two injection points: the repeated path string and the Referer header. Panel (b) illustrates a more application-specific error, a request to follow a non-existent user, resulting in a single injection point: the username of the user to be followed within the social network in the POST request.
@@ H.L. Mencken
Agents are task-driven systems designed to complete assigned objectives as efficiently as possible [7]. When attackers are aware of the general goal the agent is designed to achieve, they can exploit the inherent eagerness of the agent to solve the given task, swaying it to perform adversarially chosen actions without necessarily subverting it completely, in contrast to traditional prompt injection [39, 23, 38].
@@
This payload consists of two main components: (1) a plausible explanation behind the incident (which we call lead), (2) and a recommended remediation strategy (which we call body).
Payload’s Components. The purpose of the lead is to present a direct, contextually appropriate rationale for the error, guiding the agent toward accepting the remediation in the body. Ideally, the lead should reflect the nature of the incident the agent is tasked with investigating. For example, in Figure 5, the lead points to a potential reachability issue with a resource. Since an attacker can have some influence on the semantics of the alerts they trigger and/or the state of telemetry at the time of agent activation (see Section 3.4.2), the resulting alerts and errors induced in the system must be jointly crafted to align with the lead in order to maximize attack effectiveness.
@@
Ultimately, this strategy can be seen as a deliberate form of reward hacking [7, 17], where an adversarially constructed shortcut solutions offer a low-effort path that the agent is likely to pursue. Unlike the traditional reward hacking phenomenon, where the agent exploits flaws that naturally occur in the environment or poorly specified reward function, here the shortcut is intentionally introduced by the adversary through deliberate pollution of the environment; thus the name adversarial reward-hacking. Other examples of adversarial reward-hacking payloads are shown in Table G.1 in Appendix G.
Tailoring Payload to Application’s Context. An attacker can improve the effectiveness of an adversarial reward-hacking payload by grounding it in deployment-specific details of $\hat{t}$. In Figure 5, for instance, the payload includes contextual information such as the version of the HTTP server nginx, which can be easily obtained through simple tools like nmap. Such information requires no special access and can be gathered during a reconnaissance phase using techniques like OSINT (Open Source Intelligence), domain scanning, and service enumeration. These insights enable the attacker to craft more contextually relevant (and thus more convincing to the agent) payloads. In Appendix B, we present how this process can be automated to generate realistic adversarial reward-hacking payloads using system data such as port scan results.
More Contextual Relevance with Decorators. We observed that the effectiveness of adversarial reward-hacking payloads is improved when combining them with decorator-like strings such as ‘‘[SOLUTION] \$PAYLOAD’’ or ‘‘#HUMAN HINT: \$PAYLOAD’’. Decorators serve two main purposes: (1) they help the payload escape the syntax constraints of the telemetry in which it is injected (see Figure 7(b)), and (2) they further help to contextualize the payload as a suitable solution for the current task. According to our experiments, this approach is particularly effective against highly capable models, such as GPT-4.1 and reasoning models. Examples of decorators are reported in Figure G.2.
@@ 3.4.2 Agent Activation for Tainted Telemetry Ingestion
Once the attacker has successfully tainted the target’s telemetry, the next step is to ensure that the AIOps agent is activated so it can begin executing its RCA/incident response. Whether explicit agent activation is required depends on the specific AIOps implementation. In our methodology review, we identified three main activation settings:
@@
    The AIOps agent runs on a fixed schedule, periodically scanning for potential faults and remediating them without relying on external events. This category also includes the other cases where the agent’s activation is entirely independent of any attacker-driven events.
To manually trigger an alert in setting (1), the attacker must perform a large number of actions within the system that mimic the behavior of a legitimate fault. As with telemetry injection, an effective strategy is to leverage actions that naturally produce errors; for example, sending repeated requests to non-existent resources within the target application, or initiating security-sensitive operations such as multiple password reset attempts or failed logins. Running AIOpsDoom’s fuzzer against the application under a high workload (see Section 3.3.2) is typically sufficient to trigger alert-rule-based errors and anomaly metrics. This means that the attacker can generally perform alert triggering and telemetry poisoning within the same step.
In setting (3), the AIOps agent runs on a periodic or arbitrary schedule and does not require any explicit triggering action. However, to maximize the impact of the adversarial reward-hacking, the attacker should leave traces of anomalous behavior in the system’s telemetry; for instance, by inducing a high number of 404/500 errors in line with chosen lead of the payload. This can be achieved using the same techniques described for setting (1) as shown in Appendix C.1. In this case, fuzzing the application is usually sufficient to pollute the telemetry and lead to manipulation.
If the system supports only setting (2), the attacker can submit a ticket reporting a plausible service issue e.g.,  “page x is not reachable anymore” if possible. This setting provides the most control over the timing and content of the alert to the attacker, and it can be seen as an idealistic (for the attacker) case (1). If the attacker lacks sufficient privileges to raise tickets within the target environment, the best strategy remains the one used in setting (3): pollute the telemetry through fuzzing and wait for independent agent activation.
@@ 3.5 A Step-By-Step Attack on AIOps
To illustrate the effect of the proposed attack strategy on AIOps’s behavior, we present an end-to-end attack demonstration against a complete and realistic AIOps-supported system. This example is based on one of the setups used in our evaluation (Section 4): the SocialNet application, which features an incident response agent built using the Flash agentic framework [58] and deployed with GPT-4o. Details about the tested agents and applications are provided in Section 4.1.
We run AIOpsDoom on the application using the example payload shown in Figure 5. After the attack, we run a post-mortem analysis of the attacked system to analyze its effect on the agent.
AIOpsDoom induces the creation of multiple tainted telemetry instances. An example of a request issued by the fuzzer, along with the corresponding tainted telemetry observed on the target, is shown in Figure 7.
@@
Legitimizing Payload with Agent-Added Content. In our experiments, we observed an unusual agentic behavior. Agents often augment the (adversarially chosen) root cause with additional context in an effort to self-contextualize the incident and its remediation. For example, in the scenario above, the agent retrieves and explicitly includes the exact distribution and version of the HTTP server running on the application (i.e., “openresty/1.15.8.1rc1”); information that was not part of the payload. We provide additional examples in Table G.1 in Appendix G. This behavior is particularly concerning for security, as it lends the generated (and potentially false) root cause and remediation a heightened sense of realism and correctness, grounded in system-specific details that are not publicly available. Consequently, the adversarially injected remediation becomes more believable and less likely to raise suspicion, thereby increasing the risk that it will bypass possible manual reviews by human operators or automated assessments by LLM-based judges before the remediation implementation.
The Flash-based agent [58] used in this example incorporates a reasoning and hindsight mechanism. To provide insight into the agent’s decision-making behavior, in Figure G.4, we present its internal thoughts before and after accessing the logs tainted with the adversarial reward-hacking payload.